The grude appesl process at aunivenity equines Prat jury be structured by selecting seven individuals randomly from a pool of thirteen students and ten faculty. (a) What is the probablity of selecting a jury of all students? (b) What is the probability of selecting ajuny of al fulit)What is the probability of selecting a jury of three students and four faculty a) What is the probability of selectng a jury of all students? Round to fve decimal places as neded) Round to fve decimal places as needed) Round to fve decimal places as needed) lWhatis the probubility f sing a jury of three students and four fecuty?
Answer #1

Given,

Total no. of students =13

Total no. of faculty = 10

Total = 13+10= 23

If a jury of 7 individual is selected the total no. of exhaustive (Outcome), denoted by 'm' is 23 CT Ст

a). If we select a jury of all from students, the total no. of favourable cases, denoted by 'm' is 13Ст

So, we have ,

13 23

Required probability = m C 1716 0.0069995960.00700 = = 23d, 245157

------------------------------------------------------------------------------------------------------------------------------------

b).

If we select a jury of all from faculty, m=10 C7

m 10C7 120 n 23C p =-=- 0.00049 245157

-------------------------------------------------------------------------------------------------------------------------

c).

If we select a jury of three from students and four from faculty. m= 13C3 × 10 Ca

286 × 210 245157 13C3 ×10 C4 =0.24499 Ст

--------------------------------------------------------------------------------------------------------------------------